# Evaluating the Evolution of YOLO (You Only Look Once) Models: A Comprehensive Benchmark Study of YOLO11 and Its Predecessors

### v1

@@ III-A3 Ships/Vessels Dataset
The Ships dataset is an extensive open-source collection containing approximately 13.5k images, collected by Siddharth Sah from numerous Roboflow datasets, curated explicitly for ship detection [52]. Each image has been manually annotated with bounding boxes in the YOLO format, ensuring precise and efficient detection of ships. This dataset features a single class, ”ship,” allowing for streamlined and focused analysis. However, the relatively small size of the target objects and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations. The dataset is essential for various practical applications such as maritime safety, fisheries management, marine pollution monitoring, defense, maritime security, and more.
@@ III-B Models
III-B1 Comparative Analysis: Ultralytics vs. Original YOLO Models
@@
Ultralytics library provides researchers and programmers various YOLO models for inference, validation, training, and export. Based on the results of Table I, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7. Concerning YOLOv6, the library only supports the configuration *.yaml files without the pre-trained *.pt models.
Inference | Validation | Training |
| YOLOv1 | No | No | No | No | ||||
| YOLOv2 | No | No | No | No | ||||
| YOLOv3-u (Ultralytics) | Yes | Yes | Yes | Yes | ||||
| YOLOv4 | No | No | No | No | ||||
| YOLOv5-u (Ultralytics) | Yes | Yes | Yes | Yes | ||||
| YOLOv6 | Yes | Yes | Yes | No | ||||
| YOLOv7 | No | No | No | No | ||||
| YOLOv8 | Yes | Yes | Yes | Yes | ||||
| YOLOv9 | Yes | Yes | Yes | Yes | ||||
| YOLOv10 | Yes | Yes | Yes | Yes | ||||
| YOLO11 | Yes | Yes | Yes | Yes |
@@
III-B2 YOLOv3u
III-B3 YOLOv5u
III-B4 YOLOv8
III-B5 YOLOv9
Programmable Gradient Information (PGI) is a significant advancement introduced in YOLOv9. PGI is a method that dynamically adjusts the gradient information during training to optimize learning efficiency. By selectively focusing on the most informative gradients, PGI helps preserve crucial information that might otherwise be lost in lightweight models. This advancement ensures the model retains the essential features for accurate object detection, improving overall performance.
In addition, YOLOv9 incorporates GELAN (Gradient Enhanced Lightweight Architecture Network), a new architectural advancement designed to improve parameter utilization and computational efficiency as illustrated in Figure 5. GELAN achieves this by optimizing the computational pathways within the network, allowing for better resource management and adaptability to various applications without compromising speed or accuracy.
III-B6 YOLOv10
YOLOv10, developed by Tsinghua University researchers, builds upon previous models’ strengths with key innovations [64]. The architecture has an enhanced CSPNet (Cross Stage Partial Network) backbone for improved gradient flow and reduced computational redundancy, as seen in Figure 6. The network is structured into three main components: the backbone, the neck, and the detection head. The neck includes PAN (Path Aggregation Network) layers for effective multiscale feature fusion. PAN is designed to enhance information flow by aggregating features from different layers, enabling the network to better capture and combine details at various scales, which is crucial for detecting objects of different sizes. At the same time, the One-to-Many Head generates multiple predictions per object during training to provide rich supervisory signals and improve learning accuracy. Moreover, this version also offers five scaled versions, from nano to extra-large.
III-B7 YOLO11
YOLO11 [26] is the latest innovation in the YOLO series developed by Ultralytics, building upon the developments of its predecessors, especially YOLOv8. This iteration offers five scaled models from nano to extra large, catering to various applications. Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and oriented object detection (OBB).
Key improvements in YOLO11 include the introduction of the C2PSA (Cross-Stage Partial with Self-Attention) module, as seen in Figure 7, which combines the benefits of cross-stage partial networks with self-attention mechanisms. This enables the model to capture contextual information more effectively across multiple layers, improving object detection accuracy, especially for small and colluded objects. Additionally, in YOLO11, the C2f block has been replaced by C3k2, a custom implementation of the CSP Bottleneck that uses two convolutions, unlike YOLOv8’s use of one large convolution. This block uses a smaller kernel, retaining accuracy while improving efficiency and speed.
III-C Hardware and Software Setup
Table III showcases the libraries and packages used throughout this paper. During this experiment, 23 models were trained of 5 different YOLO versions found in Table IV. To ensure a fair comparative analysis, similar hyperparameters were used throughout the whole experiment on all models found below in table V. We have used 2 NVIDIA RTX 4090 GPUs for the training and evaluation, each with 16,384 CUDA cores.
For the traffic signs dataset, undersampling techniques were applied to ensure a balanced dataset, reducing the number of images from 4381 to 3233 images split into 70% training, 20% validation, and 10% testing. After balancing the dataset, 24 classes remained, with an average of 100 images per class in the training dataset. This dataset contains numerous images of traffic sizes of different sizes that render it suitable for diverse object detection.
The Africa Wildlife dataset contains 1504 images distributed equally among all 4 classes following the 70% training, 20% validation, and 10% testing split. This dataset is used for large object detection in this section,
The Ships dataset contains 13.4k images and is divided into 70% training, 20% validation, and 10% testing. It has only one class (ship) and is focused on small object detection.
| Name | Version | Description |
| Python | 3.12 | Programming language |
| Ubuntu | 22.04 | Linux operating system |
| CUDA | 12.5 | Platform for GPU based processing |
| cuDNN | 8.9.7 | CUDA library for deep neural networks |
| Ultralytics | 8.2.55 | YOLO Object Detection Library |
| WandB | 0.17.4 | ML experiment tracking |
@@ III-C Hardware and Software Setup
| YOLO11x (extra-large) |
@@ III-D Metrics
This study evaluates the performance of YOLO models using three primary metrics: accuracy, computational efficiency, and size. The accuracy metrics include Precision, Recall, mAP50 (Mean Average Precision at an IoU (Intersection over Union) threshold of 0.50), and mAP50-95 (Mean Average Precision across IoU (Intersection over Union) thresholds from 0.50 to 0.95). Precision [41] measures the ratio of correctly predicted observations to the total predicted observations, thus highlighting the occurrence of False Positives. Conversely, Recall [41] measures the ratio of correctly predicted observations to all actual observations, thus emphasizing the occurrence of False Negatives. Both mAP50 and mAP50-95 [70] provide a comprehensive summary of Precision and Recall. While mAP50 calculates the Mean Average Precision at an IoU threshold of 0.50, mAP50-95 extends this calculation across multiple IoU thresholds from 0.50 to 0.95, with a step size of 0.05.
Regarding computational efficiency metrics, Preprocessing Time, Inference Time, and Postprocessing Time will be utilized to evaluate the model’s speed. Preprocessing Time refers to the duration taken to prepare raw data for input into the model. Inference Time is the duration required for the model to process the input data and generate predictions. Postprocessing Time denotes the time needed to convert the model’s raw predictions into a final, usable format. These metrics were measured using a sample of images for testing after training the models. Additionally, the GFLOPs (Giga Floating-Point Operations Per Second) measure the computational power for the model training, reflecting its efficiency. In contrast, the size metric reflects the actual disk size of the model and the number of its parameters.
These metrics are essential for providing a comprehensive overview of YOLO models’ performance, allowing for effective comparison and evaluation. By employing these metrics, we can thoroughly assess the accuracy and efficiency of different YOLO model versions, ensuring a robust benchmark for their performance and application in various real-world scenarios.
@@ IV-A1 Traffic Signs Dataset
| YOLO11x | 0.819 | 0.816 | 0.885 | 0.784 | 0.9 | 6.1 | 0.4 | 7.4 | 109 | 194.8 |
Table VI presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size. The Traffic Signs dataset is a medium-sized dataset with varied object sizes, making it favorable for benchmarking. The results highlight the effectiveness of YOLO models in detecting traffic signs, demonstrating a range of precision. The highest mAP50-95 was 0.799, while the lowest recorded precision was 0.64. On the other hand, the highest mAP50 is 0.893 while the lowest is 0.722. The substantial gap between the mAP50 and mAP50-95 results suggests that the models encounter difficulties in uniformly handling traffic signs with different sizes at higher thresholds, reflecting areas for potential improvement in their detection algorithms.
@@
Figure 9 elucidates the trade-off between precision and recall taking the size of the models into consideration. Models such as YOLO11m, YOLO10l, YOLOv9m, YOLOv5ux, and YOLO11l exhibit high precision and recall, specifically with YOLO11m achieving a precision of 0.898 and a recall of 0.826 while having a size of 67.9Mb, and YOLOv10l achieving a precision of 0.873 and a recall of 0.807 with a significantly bigger size (126.8 Mb). In contrast, smaller models such as YOLOv10n (precision 0.722, recall 0.602), YOLOv8n (precision 0.749, recall 0.688), and YOLO11n (precision 0.768, recall 0.695) underperform in both metrics. This underscores the superior performance of larger models on the Traffic Signs dataset. Moreover, the high precision (0.849) and low recall (0.701) of YOLOv5um indicate a propensity for false negatives, while YOLOv3u’s high recall (0.849) and low precision (0.75) suggest a tendency for false positives.
@@
In terms of computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figures 10 and 11. YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u-tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models. However, the data indicates that YOLOv9e, YOLOv9m, YOLOv9c, and YOLOv9s are the least efficient, with inference times of 16.1ms, 12.1ms, 11.6ms, and 11.1ms, and GFLOPs count of 189.4, 76.7, 102.6, and 26.8 respectively. These findings delineate a clear trade-off between accuracy and computational efficiency.
@@
When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model. It achieves a mAP50-95 of 0.795, an inference time of 2.4ms, a model size of 38.8Mb, and a 67.9 GFLOPs count, as detailed in Figures 8, 10, 11, and Table VI. This is followed by YOLO11l (mAP50-95 of 0.794, inference time of 4.6ms, size of 49Mb, and 86.8 GFLOPs count), and YOLOv10m (mAP50-95 of 0.781, inference time of 2.4ms, size of 32.1Mb, 63.8 GFLOPs count). These results highlight the robustness of these models in detecting traffic signs of various sizes while maintaining short inference times and small model sizes. Notably, the YOLO11 and YOLOv10 families significantly outperform other YOLO families, in terms of accuracy and computational efficiency in this dataset, as their models consistently surpass counterparts from other families.
@@ IV-A2 Africa Wildlife Dataset
Total Time | Size | GFLOPs | ||||||
| YOLOv3u | 0.91 | 0.88 | 0.943 | 0.803 | 0.5 | 6.2 | 0.4 | 7.1 | 207.86 | 282.2 | ||||||
| YOLOV3u tiny | 0.897 | 0.866 | 0.921 | 0.725 | 0.7 | 0.7 | 0.4 | 1.8 | 24.44 | 19.1 | ||||||
| YOLOv5un | 0.949 | 0.862 | 0.948 | 0.791 | 1.1 | 0.5 | 0.4 | 2 | 5.65 | 7 | ||||||
| YOLOv5us | 0.924 | 0.882 | 0.853 | 0.804 | 1 | 0.8 | 0.4 | 2.2 | 18.58 | 23.8 | ||||||
| YOLOv5um | 0.935 | 0.887 | 0.947 | 0.807 | 0.6 | 2.1 | 0.4 | 3.1 | 50.54 | 64 | ||||||
| YOLOv5ul | 0.916 | 0.881 | 0.948 | 0.797 | 0.7 | 3.3 | 0.4 | 4.4 | 106.85 | 135 | ||||||
| YOLOv5ux | 0.932 | 0.867 | 0.946 | 0.797 | 0.5 | 6.6 | 0.4 | 7.5 | 195.2 | 246.2 | ||||||
| YOLOv8n | 0.932 | 0.908 | 0.953 | 0.794 | 1.1 | 0.5 | 0.4 | 2 | 6.55 | 8.2 | ||||||
| YOLOv8s | 0.962 | 0.88 | 0.943 | 0.812 | 0.9 | 1 | 0.4 | 2.3 | 22.59 | 28.7 | ||||||
| YOLOv8m | 0.928 | 0.909 | 0.953 | 0.822 | 0.8 | 2.5 | 0.4 | 3.7 | 52.12 | 78.9 | ||||||
| YOLOv8l | 0.942 | 0.898 | 0.957 | 0.817 | 0.9 | 3.9 | 0.4 | 5.2 | 87.77 | 165.1 | ||||||
| YOLOv8x | 0.912 | 0.906 | 0.953 | 0.819 | 0.5 | 7.1 | 0.4 | 8 | 136.9 | 257.6 | ||||||
| YOLOv9t | 0.944 | 0.875 | 0.948 | 0.825 | 1.3 | 1.1 | 0.4 | 2.8 | 4.93 | 7.7 | ||||||
| YOLOv9s | 0.921 | 0.897 | 0.956 | 0.832 | 1 | 1.2 | 0.4 | 2.6 | 15.33 | 26.9 | ||||||
| YOLOv9m | 0.924 | 0.901 | 0.952 | 0.823 | 0.9 | 2.8 | 0.4 | 4.1 | 40.98 | 76.5 | ||||||
| YOLOv9c | 0.934 | 0.897 | 0.96 | 0.83 | 0.9 | 3.4 | 0.4 | 4.7 | 51.8 | 102.7 | ||||||
| YOLOv9e | 0.932 | 0.864 | 0.944 | 0.809 | 0.5 | 7.6 | 0.4 | 8.5 | 117.5 | 189.3 | ||||||
| YOLOv10n | 0.901 | 0.9 | 0.936 | 0.786 | 1.1 | 0.7 | 0.2 | 2 | 5.59 | 8.2 | ||||||
| YOLOv10s | 0.929 | 0.888 | 0.947 | 0.799 | 0.9 | 1.1 | 0.2 | 2.2 | 15.9 | 24.5 | ||||||
| YOLOv10m | 0.91 | 0.88 | 0.945 | 0.8 | 1 | 2.4 | 0.2 | 3.6 | 32.1 | 63.4 | ||||||
| YOLOv10b | 0.905 | 0.899 | 0.944 | 0.809 | 0.8 | 3.2 | 0.2 | 4.2 | 39.7 | 98 | ||||||
| YOLOv10l | 0.922 | 0.894 | 0.944 | 0.8 | 0.7 | 3.8 | 0.2 | 4.7 | 50 | 126.3 | ||||||
| YOLOv10x | 0.96 | 0.862 | 0.949 | 0.819 | 0.8 | 6.3 | 0.2 | 7.3 | 61.4 | 169.8 | ||||||
| YOLO11n | 0.964 | 0.877 | 0.964 | 0.802 | 1.1 | 0.7 | 0.4 | 2.2 | 5.35 | 6.3 | ||||||
| YOLO11s | 0.952 | 0.892 | 0.959 | 0.8 | 1.1 | 1 | 0.4 | 2.5 | 18.4 | 21.3 | ||||||
| YOLO11m | 0.922 | 0.906 | 0.96 | 0.81 | 0.9 | 2.4 | 0.4 | 3.7 | 38.8 | 67.7 | ||||||
| YOLO11l | 0.937 | 0.896 | 0.965 | 0.805 | 0.9 | 3 | 0.4 | 4.3 | 49 | 86.6 | ||||||
| YOLO11x | 0.908 | 0.886 | 0.945 | 0.795 | 0.6 | 6.1 | 0.4 | 7.1 | 109 | 194.4 |
@@
YOLOv10n, YOLOv8n, and YOLOv3u-tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively. The first two models share the same processing speed and GFLOPs count, as showcased in Figures 14 and 15. Conversely, YOLOv9e exhibits the slowest processing time at 11.2ms and a GFLOPs count of 189.3, followed by YOLOv5ux at 7.5ms and 246.2 GFLOPs count. These results indicate that larger models tend to require more processing time and hardware usage compared to smaller models, emphasizing the trade-off between model size and processing efficiency.

### v2

@@ III-A3 Ships/Vessels Dataset
The Ships/Vessels dataset is an extensive open-source collection containing approximately 13.5k images, collected by Siddharth Sah from numerous Roboflow datasets, curated explicitly for ship detection [55]. Each image has been manually annotated with bounding boxes in the YOLO format, ensuring precise and efficient detection of ships. This dataset features a single class, ”ship,” and is divided into 70% training, 20% validation, and 10% testing allowing for streamlined and focused analysis. Notably, no data augmentation techniques were applied during the experiment, which ensures that the model’s performance is evaluated on the raw, unaltered dataset. However, the relatively small size of the target objects and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations. The dataset is essential for various practical applications such as maritime safety, fisheries management, marine pollution monitoring, defense, maritime security, and more.
@@ III-B Models
III-B1 Ultralytics vs. Original YOLO
@@
Ultralytics library provides researchers and programmers with various YOLO models for inference, validation, training, and export. Based on the results of Table I, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7. Concerning YOLOv6, the library only supports the configuration *.yaml files without the pre-trained *.pt models.
Inference | Validation | Training | ||
| YOLOv1 | No | No | No | ||
| YOLOv2 | No | No | No | ||
| YOLOv3u | Yes | Yes | Yes | ||
| YOLOv4 | No | No | No | ||
| YOLOv5u | Yes | Yes | Yes | ||
| YOLOv6 | Yes | Yes | Yes | ||
| YOLOv7 | No | No | No | ||
| YOLOv8 | Yes | Yes | Yes | ||
| YOLOv9 | Yes | Yes | Yes | ||
| YOLOv10 | Yes | Yes | Yes | ||
| YOLO11 | Yes | Yes | Yes | ||
| YOLOv12 | Yes | Yes | Yes |
@@
In total, 33 models from 7 different YOLO versions were trained on 3 different datasets as seen in table III. Each YOLO version includes several scaled models, (e.g., YOLOv5u, YOLOv5un, YOLOv5us, and YOLOv5ux), with the suffixes denoting model size and complexity, such as ”n” for nano, ”s” for small, ”m” for medium, ”l” for large, ”x” for extra-large, ”t” for tiny, ”c” for compact, ”b” for balanced, and ”e” for extended. These models offer a variety of trade-offs between accuracy and inference speed, as discussed in the following sections.
@@
| YOLO11x (extra-large) | |
| YOLOv12 | YOLO11n (nano) |
| YOLOv12s (small) | |
| YOLOv12m (medium) | |
| YOLOv12l (large) | |
| YOLOv12x (extra-large) |
III-B2 YOLOv3u
III-B3 YOLOv5u
III-B4 YOLOv8
III-B5 YOLOv9
Programmable Gradient Information (PGI) is a significant advancement introduced in YOLOv9. PGI is a method that dynamically adjusts the gradient information during training to optimize learning efficiency. By selectively focusing on the most informative gradients, PGI helps preserve crucial information that might be lost in lightweight models. This advancement ensures the model retains the essential features for accurate object detection, improving overall performance.
In addition, YOLOv9 incorporates Gradient Enhanced Lightweight Architecture Network (GELAN), a new architectural advancement designed to improve parameter utilization and computational efficiency as illustrated in Figure 7. GELAN achieves this by optimizing the computational pathways within the network, allowing for better resource management and adaptability to various applications without compromising speed or accuracy.
III-B6 YOLOv10
YOLOv10, developed by Tsinghua University researchers, builds upon previous models’ strengths with key innovations [68]. The architecture has an enhanced Cross Stage Partial Network (CSPNet) backbone for improved gradient flow and reduced computational redundancy, as seen in Figure 8. The network is structured into three main components: the backbone, the neck, and the detection head. The neck includes Path Aggregation Network (PAN) layers for effective multiscale feature fusion. PAN is designed to enhance information flow by aggregating features from different layers, enabling the network to better capture and combine details at various scales, which is crucial for detecting objects of different sizes. At the same time, the One-to-Many Head generates multiple predictions per object during training to provide rich supervisory signals and improve learning accuracy. Five scaled versions are available, from nano to extra-large.
III-B7 YOLO11
YOLO11 [28] is one of the latest additions to the YOLO series developed by Ultralytics, building upon the developments of its predecessors, especially YOLOv8. This iteration offers five scaled models from nano to extra large, catering to various applications. Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and oriented object detection (OBB).
Key improvements in YOLO11 include the introduction of the Cross-Stage Partial with Self-Attention (C2PSA) module, as seen in Figure 9, which combines the benefits of cross-stage partial networks with self-attention mechanisms. This enables the model to capture contextual information more effectively across multiple layers, improving object detection accuracy, especially for small and colluded objects. Additionally, in YOLO11, the C2f block has been replaced by C3k2, a custom implementation of the CSP Bottleneck that uses two convolutions, unlike YOLOv8’s use of one large convolution. This block uses a smaller kernel, retaining accuracy while improving efficiency and speed.
III-B8 YOLOv12
YOLOv12 [63] is the latest evolution in the YOLO series, introducing an attention-centric design that significantly enhances both speed and accuracy. It continues the trend of offering five scalable models (Nano to Extra Large), making it adaptable for a wide range of applications such as object detection, instance segmentation, and oriented object detection (OBB). As seen in Figure 10, YOLOv12 introduces the Area Attention (A2) module, which maintains a large receptive field while drastically reducing computational complexity, allowing the model to enhance speed without compromising accuracy. Additionally, it features Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
Moreover, YOLOv12 uses FlashAttention to minimize memory access overhead, closing the speed gap with CNNs. It adjusts the MLP ratio from 4 to 1.2, enhancing runtime efficiency, and removes positional encodings for a cleaner, faster model without losing detection accuracy. YOLOv12 also reduces complexity by using a single R-ELAN block in the last stage of the backbone instead of stacking three attention/CNN blocks. It replaces linear layers with convolutional layers and batch normalization, maximizing computational efficiency and achieving state-of-the-art latency-accuracy trade-offs.
III-C Hardware and Software Setup
Table IV presents the libraries and software versions used throughout this study. The experiments were conducted with Python 3.12, Ubuntu 22.04, CUDA 12.5, and cuDNN 8.9.7 for GPU acceleration. Ultralytics 8.2.55 was used for model training, while WandB 0.17.4 was utilized for tracking experiments. To ensure a fair comparison, similar hyperparameters were used across all models, as outlined in Table V. The key hyperparameters include 100 epochs, the AdamW optimizer, a batch size of 16, and an image size of (640, 640). The learning rate was initially set to 0.0001 and increased to 0.01, while the dropout rate was set to 0.15. The dataset was split into 70% for training, 20% for validation, and 10% for testing with no data augmentation techniques applied to any dataset. The experiments were carried out using 2 NVIDIA RTX 4090 GPUs, each with 16,384 CUDA cores.
| Name | Version |
| Python | 3.12 |
| Ubuntu | 22.04 |
| CUDA | 12.5 |
| cuDNN | 8.9.7 |
| Ultralytics | 8.2.55 |
| WandB | 0.17.4 |
@@ III-D Metrics
This study evaluates the performance of YOLO models using three primary metrics: accuracy, computational efficiency, and size. The accuracy metrics include Precision, Recall, Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50 (mAP50), and Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50-0.95 (mAP50-95). Precision [44] measures the ratio of correctly predicted observations to the total predicted observations, thus highlighting the occurrence of False Positives. Conversely, Recall [44] measures the ratio of correctly predicted observations to all actual observations, thus emphasizing the occurrence of False Negatives. Both mAP50 and mAP50-95 [75] provide a comprehensive summary of Precision and Recall. While mAP50 calculates the Mean Average Precision at an IoU threshold of 0.50, mAP50-95 extends this calculation across multiple IoU thresholds from 0.50 to 0.95, with 0.05 step size.
Regarding computational efficiency metrics, Preprocessing Time, Inference Time, and Postprocessing Time will be utilized to evaluate the model’s speed. Preprocessing Time refers to the duration taken to prepare raw data for input into the model. Inference Time is the duration required for the model to process the input data and generate predictions. Postprocessing Time denotes the time needed to convert the model’s raw predictions into a final, usable format. These metrics were measured using a sample of images for testing after training the models. Additionally, the Giga Floating-Point Operations Per Second (GFLOPs) measure the computational power for the model training, reflecting its efficiency. In contrast, the size metric reflects the actual disk size of the model and the number of its parameters.
These metrics are essential for providing a comprehensive overview of YOLO models’ performance, allowing for effective comparison and evaluation. These metrics allows a thorough assessment of the accuracy and efficiency of different YOLO models, ensuring a robust benchmark for their performance and applications in various real-world scenarios.
@@ IV-A1 Traffic Signs Dataset
| YOLO11x | 0.819 | 0.816 | 0.885 | 0.784 | 0.9 | 6.1 | 0.4 | 7.4 | 109 | 194.8 | ||||||||
| YOLOv12n | 0.883 | 0.684 | 0.802 | 0.708 | 1.2 | 4.6 | 0.4 | 6.2 | 5.34 | 6.4 | ||||||||
| YOLOv12s | 0.761 | 0.839 | 0.877 | 0.78 | 1.6 | 4.3 | 0.5 | 6.4 | 18.1 | 21.4 | ||||||||
| YOLOv12m | 0.868 | 0.811 | 0.89 | 0.79 | 0.9 | 4.9 | 0.4 | 6.2 | 39 | 67.3 | ||||||||
| YOLOv12l | 0.826 | 0.836 | 0.877 | 0.781 | 0.9 | 5.5 | 0.4 | 6.8 | 51.2 | 89.6 | ||||||||
| YOLOv12x | 0.832 | 0.849 | 0.889 | 0.793 | 1 | 9.5 | 0.5 | 11 | 114 | 198.9 |
Table VI presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size. The Traffic Signs dataset is a medium-sized dataset with varied object sizes, making it favorable for benchmarking. The results highlight the effectiveness of YOLO models in detecting traffic signs, demonstrating a range of precision. The highest mAP50-95 was 0.799, while the lowest recorded precision was 0.64. On the other hand, the highest mAP50 is 0.893 while the lowest is 0.722. The substantial gap between the mAP50 and mAP50-95 results suggests that the models have difficulties in uniformly handling traffic signs with different sizes at higher thresholds, reflecting areas for potential improvement in their detection algorithms.
@@
Figure 12 elucidates the trade-off between precision and recall taking the size of the models into consideration. Models such as YOLO11m, YOLO10l, YOLOv12m, YOLOv9m, YOLOv5ux, and YOLO11l exhibit high precision and recall, specifically with YOLO11m achieving a precision of 0.898 and a recall of 0.826 while having a size of 67.9Mb, and YOLOv12m achieving a precision of 0.868 and a recall of 0.811 with a significantly smaller size (39 Mb). In contrast, smaller models such as YOLOv10n (precision 0.722, recall 0.602), YOLOv8n (precision 0.749, recall 0.688), and YOLO11n (precision 0.768, recall 0.695) underperform in both metrics. This underscores the superior performance of larger models on the Traffic Signs dataset. Moreover, the high precision (0.883) and low recall (0.684) of YOLOv12n indicate a propensity for false negatives, while YOLOv3u’s high recall (0.849) and low precision (0.75) suggest a tendency for false positives.
@@
In terms of computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figure 13. YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u-tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models. However, the data indicates that YOLOv9e, YOLOv9m, YOLOv9c, and YOLOv9s are the least efficient, with inference times of 16.1ms, 12.1ms, 11.6ms, and 11.1ms, and GFLOPs count of 189.4, 76.7, 102.6, and 26.8, respectively. These findings delineate a clear trade-off between accuracy and computational efficiency.
@@
When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model. It achieves a mAP50-95 of 0.795, an inference time of 2.4ms, a model size of 38.8Mb, and a 67.9 GFLOPs count, as detailed in Figures 11, 13, and Table VI. This is followed by YOLO11l (mAP50-95 of 0.794, inference time of 4.6ms, size of 49Mb, and 86.8 GFLOPs count), and YOLOv10m (mAP50-95 of 0.781, inference time of 2.4ms, size of 32.1Mb, 63.8 GFLOPs count). These results highlight the robustness of these models in detecting traffic signs of various sizes while maintaining short inference times and small model sizes.
On the contrary, YOLOv9e demonstrates poor performance overall, achieving a relatively low mAP50-95 of 0.764, an inference time of 16.1ms, a large model size of 117.5MB, and one of the highest GFLOPs counts at 198.4. Additionally, YOLOv5um, YOLOv8m, and YOLOv8s performed suboptimally, further indicating the inconsistency in performance among earlier model families.
Notably, the YOLO11 and YOLOv12 families significantly outperform other YOLO families in terms of accuracy and computational efficiency. Their models consistently surpass counterparts from the YOLOv3u, YOLOv5u, YOLOv8, YOLOv9, YOLOv10 families, demonstrating their ability to balance precision and speed effectively in the detection of objects with varying sizes.
@@ IV-A2 Africa Wildlife Dataset
GFLOPs | Size | ||||||||
| YOLO11x | 0.819 | 0.816 | 0.885 | 0.784 | 0.9 | 6.1 | 0.4 | 7.4 | 109 | 194.8 | ||||||||
| YOLOv12n | 0.883 | 0.684 | 0.802 | 0.708 | 1.2 | 4.6 | 0.4 | 6.2 | 5.34 | 6.4 | ||||||||
| YOLOv12s | 0.761 | 0.839 | 0.877 | 0.78 | 1.6 | 4.3 | 0.5 | 6.4 | 18.1 | 21.4 | ||||||||
| YOLOv12m | 0.868 | 0.811 | 0.89 | 0.79 | 0.9 | 4.9 | 0.4 | 6.2 | 39 | 67.3 | ||||||||
| YOLOv12l | 0.826 | 0.836 | 0.877 | 0.781 | 0.9 | 5.5 | 0.4 | 6.8 | 51.2 | 89.6 | ||||||||
| YOLOv12x | 0.832 | 0.849 | 0.889 | 0.793 | 1 | 9.5 | 0.5 | 11 | 114 | 198.9 |
@@
YOLOv10n, YOLOv8n, and YOLOv3u-tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively. The first two models share the same processing speed and GFLOPs count, as showcased in Figure 16. Conversely, YOLOv12x exhibits the slowest processing time at 12.8ms and a GFLOPs count of 198.5, followed by YOLOv9e at 8.5ms and 189.3 GFLOPs count. These results indicate that larger models tend to require more processing time and hardware usage compared to smaller models, emphasizing the trade-off between model size and processing efficiency.